Calculate the pH at the equivalence point in titrating 0.051 M solutions of each of the following with 0.051 M NaOH.

(a) perbromic acid (HBrO4)

pH =

(b) carbonic acid (H2CO3), Ka = 4.3e-07

pH =

(c) boric acid (H3BO3), Ka = 5.8e-10

pH =
Answer #1

1)HBrO, is the strong acid. at equival ence point pH-7. 0.051x0.051 0.051+0.051 = 0.0255 MMcid 0.0255 = 2.3×10-8 = 5.86× 10.1